OpenStudy (anonymous):

If 1/4cotx - 11 = -9, then cotx = _____. ans:1/2,5,8, or 80

OpenStudy (rulnick):

cot x = 8, assuming by 1/4cotx you mean (1/4)cot x.

OpenStudy (anonymous):

yes i meant (1/4)cotx

OpenStudy (ash2326):

We have \[ \frac{1}{4} \cot x -11=-9\] Add 11 to both sides \[\frac{1}{4} \cot x -11+11=-9+11\] we get \[\frac{1}{4} \cot x=2\] multiply 4 to both sides \[\cot x=8\]